A textbook store sold a combined total of 200
psychology and chemistry textbooks in a week. The number of chemistry textbooks sold was 76
less than the number of psychology textbooks sold. How many textbooks of each type were sold?

Answer by ikleyn(52813)

.
C = P - 76


P + (P-76) = 200


2P = 200 + 76 = 276

P = 276/2 = 138     psychology books sold.  ANSWER


C = 200 - 138 = 62  chemistry book sold.    ANSWER


CHECK.   138 - 62 = 76.    ! Correct !

Solved.